OpenStudy (anonymous):

Solve this equation. 4 + 18 + c = 25

OpenStudy (anonymous):

what is 4+18?

OpenStudy (anonymous):

@ashxoxoley are you there?

OpenStudy (anonymous):

22

OpenStudy (squirrels):

22+c=25 is what you have now Now you must subtract 22 by itself to isolate c and subtract 22 to the other side

OpenStudy (squirrels):

So it will look like this 22+c=25 -22 -22 ----------- ?????????????? <---Just answer that and c will be your answer